OpenStudy (beccab003):

Solve the following system of equations. x^2 + y^2 = 9 x^2 - 3y = 9

OpenStudy (beccab003):

These are the options: (-3, 0), (3, 0),(-3, 9),(3, 9) (-3, 9),(-3, 0) (-3, 0), (0, -3), (3, 0) (-3, 9) I am very confused on how to do this problem because there is a 3y and y^2

OpenStudy (anonymous):

well since they both = 9 its safe to say they equal each other right @BeccaB003

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

well since they both = 9 its safe to say they equal each other right @BeccaB003

OpenStudy (beccab003):

I am here. Yes, I suppose they would equal each other.

OpenStudy (welshfella):

x^2 + y^2 = x^2 - 3y

OpenStudy (anonymous):

so x^2 + y^2 = x^2 - 3y now solve do -x^2 to get y^2+-3y

OpenStudy (anonymous):

i mean y^2=-3y

OpenStudy (welshfella):

subtract x^2 from both sides y^2 = - 3y

OpenStudy (beccab003):

@welshfella @DaBest21 But how do you solve with that problem if there is y and y^2?

OpenStudy (welshfella):

can you solve that?

OpenStudy (beccab003):

I thought you used substitution or elimination for problems like this.

OpenStudy (beccab003):

I don't understand how you would solve for y^2=3y ---they are two different types of variables. you can't combine them or anything.

OpenStudy (welshfella):

well i have used elimination - i eliminated x^2

OpenStudy (welshfella):

y^2 + 3y = 0 y(y + 3) = 0 y = ?

OpenStudy (welshfella):

you factor them as above

OpenStudy (beccab003):

y(y+3)=0 <---------- is factored

OpenStudy (beccab003):

I don't know how to simplify that more.

OpenStudy (welshfella):

when 2 functions are multiplied together and the result is 0 then either one or the other is zero. I other words there are 2 solutions y(y + 3) = 0 y = 0 or y + 3 = 0 y = 0 y= -3 0 and -3 are the 2 solutions

OpenStudy (beccab003):

So then you plug them both into the original equation to find x?

OpenStudy (welshfella):

yes plug them into x^2 - 3y = 9

OpenStudy (welshfella):

so when y = -3 x^2 -3(-3) = 9 x^2 + 9 = 9 x^2 = 9-9 = 0 x = 0 so one solution is (0,-3)

OpenStudy (beccab003):

Okay so, x^2 -3(0) = 9 equals: x^2 = 9 = x = 3

OpenStudy (beccab003):

So another solution is (3,0)

OpenStudy (welshfella):

not quite sqrt x^2 is also -3

OpenStudy (welshfella):

right (3,0) is one solution and a third one is (-3,0)

OpenStudy (welshfella):

* square root of 9 is also -3

OpenStudy (beccab003):

How can the square root of 9 be -3?

OpenStudy (welshfella):

so the 3rd option is the correct one.

OpenStudy (anonymous):

c

OpenStudy (welshfella):

3 * 3 = 9 also -3 * -3 = 9

OpenStudy (beccab003):

Okay, Thank you so much for your help! :)

OpenStudy (welshfella):

yw